# Robust iterative methods for linear systems with saddle point structure

### v1

@@ 1 Introduction
where $A\in\mathbb{R}^{n\times n}$, $B,C\in\mathbb{R}^{n\times m}$ with $m\leq n$ and $\mathop{\rm rank}\nolimits B=\mathop{\rm rank}\nolimits C=k\leq m$. Furthermore, we assume that $A+A^{T}$ is positive semidefinite (denoted as $A+A^{T}\geq 0$). In our theoretical analysis, we allow that $k<m$ which means that the system is singular. In the presented numerical methods, however, we assume that the linear system has a unique solution, which implies that $B$ and $C$ are of rank $k=m$, since this is the property that holds in almost all of the applications. However, we will also discuss the treatment of the case $k<m$ when this is the case.
@@
In most applications, one has $C=B$ but the case $B\neq C$ also arises in practice, see the examples presented in Section 4. Very often the two components of the solution vector represent different quantities, such as e.g. a state vector $x$ and a Lagrange multiplier $y$, that penalizes the constraints. It is then important to make sure that the constraints, i.e., the second equation in (1) is satisfied accurately.
@@
A classical purely algebraic approach to the solution or the construction of preconditioners for (1) is based on the (approximate) block $LU$-factorization of the coefficient matrix, see [14], that leads to a well-known scheme which then requires the approximate iterative solution of a smaller linear system where the coefficient matrix is an approximation of the Schur complement $S=-C^{T}A^{-1}B$ which is, however, typically a dense matrix. Another common technique is the nullspace approach, see [30], where one projects the problem onto a subproblem that is associated to the nullspace of $C^{T}$, which is often easily identified in practice. In the complement of this nullspace, the solution is fixed by the second equation in (1), so that the problem can be reduced to the nullspace. Note that the nullspace method has recently been extended to the case of a nonzero $(2,2)$ block, see [30].
As a modification to these well-known approaches of the nullspace method, our new method is an inner-outer iteration scheme that is based on an approximation of the classical nullspace method, see [30], and the recently introduced shifted skew-symmetric preconditioning [22]. It employs iterative schemes for the occurring least squares problems, combined with classical projection type iterative methods, see e.g. [2, 29, 31], as outer iteration. The method will be particularly interesting in the context of implicit discretization methods for linear time-varying or nonlinear dissipative Hamiltonian differential equations, where a sequence of slightly varying systems has to be solved; see [16, 23].
Let us recall the nullspace method for the general class. Given a linear system,
$W\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}A&B\\ -C^{T}&0\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}=\begin{bmatrix}f\\ g\end{bmatrix},$ (2)
suppose that we can easily compute matrices $Z$ and $U\in\mathbb{R}^{n\times(n-m)}$ whose columns form bases for the nullspaces of $B^{T}$ and $C^{T}$, respectively (i.e., $B^{T}Z=0$ and $C^{T}U=0$).
The second block equation is an underdetermined linear system $-C^{T}x=g$ which may have infinitely many solutions. Suppose that a particular solution $\hat{x}$ is given such that $-C^{T}\hat{x}=g$. Then, with $x=\hat{x}+\bar{x}$, we can solve the equivalent system
@@
$By=f-Ax.$ (6)
Note that to solve the over- and under-determined systems we may use a linear least squares approach via an approximate $QR$-decomposition. The basic nullspace method is summarized in Algorithm 1.
1:$A\in\mathbb{R}^{n\times n}$, $B,C\in\mathbb{R}^{n\times m}$, $f\in\mathbb{R}^{n}$ and $g\in\mathbb{R}^{m}$ with $m<n$.
2:$x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$
@@
4:Find a particular solution, $\hat{x}\in\mathbb{R}^{n}$, of $-C^{T}x=g$.
5:Solve $Z^{T}AUz=Z^{T}(f-A\hat{x})$.
@@
Note that in the symmetric and generalized case, where $C=B$, the method simplifies since one needs to compute only one nullspace basis $Z$, and lines 3 and 4 change to $Z^{T}AZz=Z^{T}(f-A\hat{x})$ and $x=\hat{x}+Zz$, respectively. We call $\mathcal{N}=Z^{T}AZ$ and $\mathcal{N}=Z^{T}AU$ the projected nullspace matrices, for the symmetric/generalized and general cases, respectively. In Section 2 we analyze the properties of the symmetric part of the projected (approximate) nullspace matrix, i.e., $\mathcal{N}^{s}=\frac{Z^{T}AU+(Z^{T}AU)^{T}}{2}$ and show under which conditions it is positive definite.
The nullspace method is exact only if the nullspace bases are determined exactly, and the reduced under- as well as over-determined systems are solved exactly. Since we cannot expect this to happen, we propose to compute all of these solutions approximately and use an outer iterative projection based Krylov subspace method to correct the approximations. In this way, the approximate general nullspace method becomes a method to apply a preconditioner. This approximate approach also deals with a major disadvantage of the exact nullspace method, that the matrices $Z$ and $U$ are usually not sparse, even if $B$ and $C$ are sparse. We can therefore employ techniques for computing sparse approximate nullspace bases, see [30] for a detailed survey of the available methods.
The remainder of the paper is organized as follows. In Section 2, we present the underlying theoretical foundations of the new schemes. Then, in Section 3 we describe the new schemes. The description of the test problems, the experimental setup and the results are given in Sections 4, 5, and 6, respectively. We conclude and discuss future work in Section 7.
@@ 2 Theoretical Analysis
Consider $A=J+R\in\mathbb{R}^{n\times n}$, where $0\neq R=R^{T}$ and $0\neq J=-J^{T}$. Then there exist a real orthogonal matrix $Q\in\mathbb{R}^{n\times n}$, and integers $n_{1}\geq n_{2}\geq\cdots\geq n_{r-1}>0$ and $n_{r},n_{r+1}\geq 0$, such that
@@
[…]
@@
The proof follows from the staircase form of [1] by a further full rank decomposition of the bottom right block.
Following the notation in [1], a matrix that can be transformed to the form (7) with $n_{r}=n_{r+1}=0$ is called hypocoercive.
For a system transformed to the form (7), one can perform a block $LDU$ factorization, see [16].
@@
Then it follows that $\hat{C}_{11}^{T}\hat{B}_{11}=K_{1}$ is nonsingular and hence $\hat{C}_{11},\hat{B}_{11}$ are nonsingular, as well as $\hat{B}_{12}=0$, $\hat{C}_{12}=0$. Then the assumption that $\mathop{\rm rank}\nolimits C=k$ implies that $\hat{B}_{22}=0$. Let
@@
Note that this transformation is an orthogonal congruence on the $A$ block and hence the property that $A+A^{T}\geq 0$ is preserved under the transformation. Note further that if $B=C$ then $Q_{2}=V_{2}$ and thus $Q=V$ and $B_{21},B_{31},B_{41}=0$.
@@
ii) For the generalized case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B=C$, and $\mathcal{W}$ is invertible, then there exists a real orthogonal matrix $Q$ such that $Q^{T}Z^{T}AZQ=\mbox{\rm diag}(A_{22},A_{33})$, with $A_{33}$ skew-symmetric and invertible. If the matrix $A$ is hypocoercive then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (8), where the last block is void and the symmetric part is positive definite.
iii) For the general case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B\neq C$, if the matrix $A$ is hypocoercive, then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (8), where the last block is void and the symmetric part is positive definite.
Note that in our algorithmic approaches, the condition in Corollary 1 that $A+A^{T}\geq 0$ can be relaxed, that it holds only in the nullspace $Z$.
The construction of the projection matrices $Z,U$ with orthonormal columns will typically lead to full matrices. So for the practical implementation we will use sparse approximate versions of the associated decompositions. In this way we will not obtain exact bases for the nullspaces, but sparse approximations $\tilde{Z},\tilde{U}$. For the symmetric and generalized case, the following result shows that if the error $\|Z-\tilde{Z}\|$ can be bounded by a a small quantity, then the approximate nullpace matrix $\tilde{Z}^{T}A\tilde{Z}$ will still have a positive definite symmetric part.
@@
Under the assumptions of Corollary 1, let in the symmetric and generalized case $\tilde{Z}$ be a (sparse) approximation of $Z$ (with columns that are not exactly orthonormal) such that $\|\tilde{Z}-Z\|_{2}\leq\epsilon\|Z\|_{2}=\epsilon$, with $\epsilon$ sufficiently small. Then $\tilde{Z}^{T}A\tilde{Z}$ still has a positive definite symmetric part.
@@ 3 A new multi-layer iterative scheme
where $V\in\mathbb{R}^{n\times n}$ is invertible and $L\in\mathbb{R}^{k\times n}$ is a lower trapezoidal matrix. Then, the rightmost $n-k$ columns of $V$ (i.e., $Z=V(:,k+1:n)$ in Matlab notation) form a basis for the nullspace. Since in our case, we are not interested in $L$, it is not stored. However, since $B$ is sparse, pivoting is usually required to avoid division by zero or small values in modulus. The construction for the nullspace of $C$ is analogous.
A pseudocode for the sparse approximate right oblique conjugation with pivoting to obtain a sparse approximate nullspace basis is presented in Algorithm 2. Hereafter, we refer to this procedure as SAROC. It produces a sparse basis since we assume that computed coefficients that are smaller than a given threshold and entries of the matrix for which we compute the nullspace that are lower than a given tolerance in absolute value are set to zero, as shown in lines 8 and 9, respectively. As a consequence, we obtain approximate projected nullspace matrices, $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{Z}$ and $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{U}$, respectively, for the symmetric/generalized and general case.
1:$B\in\mathbb{R}^{n\times m}$ with $\mathop{\rm rank}\nolimits(B^{T})=k\leq m$, $\rho$ (threshold) and $\tau$ (drop tolerance) $\triangleright$ $b_{i}$: is the $i^{th}$ column of $B$
2:$\widetilde{Z}\in\mathbb{R}^{n\times(n-k)}$triangleright $z_{i}$: is the $i^{th}$ column of $\widetilde{Z}$
3:$v_{i}=e_{i}$ for $(1\leq i\leq n)$
4:for $i=1,\ldots,m$ do
5:     $\sigma=b_{i}^{T}[v_{i},v_{i+1},\ldots,v_{n}]$triangleright compute coefficient vector $\sigma\in\mathbb{R}^{n-i+1}$
@@
7:     $v_{i}\leftrightarrow v_{i-l^{*}+1}$triangleright Swap the corresponding columns of $V$
8:     $\sigma_{1}\leftrightarrow\sigma_{l^{*}}$triangleright Swap the corresponding coefficients
@@
Rather than computing the classical Cholesky factorization of the symmetric part $\widetilde{\mathcal{N}}^{s}$ of the approximate projected nullspace matrix, to avoid explicitly forming this matrix, we opt for a factorized approximate inverse based Cholesky factorization. If $\widetilde{\mathcal{N}}^{s}$ is not positive definite, then the Cholesky factorization will detect this. If $\widetilde{\mathcal{N}}^{s}$ is indefinite (with a few negative eigenvalues), it would still be possible to obtain a shifted-skew symmetric system, as shown in [22], but with an additional cost.
We obtain the FSAI of $\widetilde{\mathcal{N}}^{s}$ using the method in [2]. Our implementation, presented in Algorithm 3, is implicit in the sense that $\widetilde{\mathcal{N}}^{s}$ is not formed explicitly, since it can be relatively dense even if $\widetilde{Z}$, $\widetilde{U}$, and $A$ are sparse. Instead, its columns are extracted, as needed in line 3. This can be easily achieved by implicit matrix-vector multiplications, $n^{s}_{i}=\widetilde{\mathcal{N}}^{s}e_{i}$. Similar to the sparse approximate right oblique conjugation, we replace computed coefficients that are smaller than a given threshold and entries, of the matrix that is factorized, that are smaller than a given tolerance in absolute value by zero, see lines 5 and 6, respectively. FSAI produces $\widetilde{W}\in\mathbb{R}^{(n-k)\times(n-k)}$ such that $\widetilde{W}^{T}\widetilde{\mathcal{N}}^{s}\widetilde{W}\approx I$. Furthermore, $\widetilde{\mathcal{N}}^{s}$ and also the skew-symmetric part $\widetilde{\mathcal{N}}^{j}$, or any other matrices resulting from further projections on them, do not need to be formed explicitly, since matrix-vector multiplications with them can be computed implicitly.
@@
1:$\widetilde{\mathcal{N}}^{s}\in\mathbb{R}^{(n-k)\times(n-k)}$, $\rho$ (threshold) and $\tau$ (drop tolerance)
2:$\widetilde{W}\in\mathbb{R}^{n\times(n-k)}$
3:$w_{i}=e_{i}$ for $(1\leq i\leq n-k)$
4:for $k=1,\ldots,n-k$ do
5:     $[\sigma_{k},\sigma_{k+1},\ldots,\sigma_{n-k}]^{T}=[w_{k},w_{k+1},\ldots,w_{n-k}]^{T}n^{s}_{i}$
$\triangleright$ $n^{s}_{i}$ is the $i^{th}$ column of $\widetilde{\mathcal{N}}^{s}$ and $w_{i}$ is the $i^{th}$ column of $\widetilde{W}$
6:     for $i=k+1,\ldots,n-k$ do
7:         if $|\sigma_{j-i+1}/\sigma_{1}|>\rho$ then
8:              $w_{i}\xleftarrow{\text{$|w_{i}(\cdot)|\geq\tau||w_{i}||_{2}$}}w_{i}-(\sigma_{i}/\sigma_{k})w_{k}$
@@
12:$\sigma_{i}=1/\sqrt{\sigma_{i}}$ for $(1\leq i\leq n-k)$
13:$\widetilde{W}=\widetilde{W}P$ where $P$ is a diagonal matrix with entries $P_{(i,i)}=\sigma_{i}$

### v2

@@ 1 Introduction
where $A\in\mathbb{R}^{n\times n}$, $B,C\in\mathbb{R}^{n\times m}$ with $m\leq n$ and $\mathop{\rm rank}\nolimits B=\mathop{\rm rank}\nolimits C=k\leq m$. Furthermore, we assume that $A+A^{T}$ is positive semidefinite (denoted as $A+A^{T}\geq 0$). In our theoretical analysis, we allow that $k<m$, which means that the system is singular. In the presented numerical methods, however, we assume that the linear system has a unique solution, which implies that $B$ and $C$ are of rank $k=m$, since this is the property that holds in almost all of the applications. We refer to [3] for the necessary and sufficient conditions for the nonsingularity of $\mathcal{W}$ under various assumptions. However, we will also discuss the treatment of the case $k<m$.
@@
In most applications, one has $C=B$ but the case $B\neq C$ also arises in practice, see the examples presented in Section 4. Very often the two components of the solution vector represent different quantities, such as e.g. a state vector $x$ and a Lagrange multiplier $y$, that penalizes the constraints.
@@
A classical purely algebraic approach to the solution or the construction of preconditioners for (1) is based on the (approximate) block $LU$-factorization of the coefficient matrix, see [15], that leads to a well-known scheme which then requires the approximate iterative solution of a smaller linear system where the coefficient matrix is an approximation of the Schur complement $S=C^{T}A^{-1}B$ which is, however, typically a dense matrix. Another common technique is the nullspace approach, see [32], where one projects the problem onto a subproblem that is associated to the nullspace of $C^{T}$ which is often easily identified in practice. While, in general for large-scale problems, computing the exact nullspace basis is very difficult [9], for some applications, it can be easily identified (see for example the list of applications and references given in [3]). In the complement of this nullspace, the solution is fixed by the second equation in (1), so that the problem can be reduced to the nullspace. Note that the nullspace method has recently been extended to the case of a nonzero $(2,2)$ block, see [32].
As a modification to these well-known approaches of the nullspace method, our new method is an inner-outer iteration scheme that is based on an approximation of the classical nullspace method, see [32], and the recently introduced shifted skew-symmetric preconditioning [23]. It employs iterative schemes for the occurring least squares problems, combined with classical projection type iterative methods, see e.g. [2, 31, 34], as an outer iteration. The method will be particularly interesting in the context of implicit discretization methods for linear time-varying or nonlinear dissipative Hamiltonian differential equations, where a sequence of slightly varying systems has to be solved; see [17, 24].
Given a linear system of the form (1), suppose that we can easily compute matrices $Z$ and $U\in\mathbb{R}^{n\times(n-m)}$ whose columns form (approximate) bases for the nullspaces of $B^{T}$ and $C^{T}$, respectively (i.e., $B^{T}Z=0$ and $C^{T}U=0$).
The second block equation in (1) is an underdetermined linear system $C^{T}x=g$ which may have infinitely many solutions. Suppose that a particular solution $\hat{x}$ is given such that $C^{T}\hat{x}=g$. Then, with $x=\hat{x}+\bar{x}$, we can solve the equivalent system
@@
By=fAx.By=f-Ax. (5)
Note that to solve the over- and under-determined systems , in general, we may use a linear least squares approach via an approximate or exact sparse $QR$ decomposition , or iteratively via a Krylov subspace method. The basic nullspace method is summarized in Algorithm 1.
[…]
@@
4:Find a particular solution, $\hat{x}\in\mathbb{R}^{n}$, of $C^{T}x=g$.
[…]
@@
[…]
The nullspace method is exact only if the nullspace bases are determined exactly, and the reduced under- as well as over-determined systems are solved exactly. Since we cannot expect this to happen, we propose to compute all of these solutions approximately and use an outer iterative projection based Krylov subspace method to correct the approximations. In this way, the approximate general nullspace method becomes a method to apply a preconditioner. This approximate approach also addresses a major disadvantage of the exact nullspace method, that the matrices $Z$ and $U$ are usually not sparse, even if $B$ and $C$ are sparse. In the approximate nullspace approach, we can employ sparse approximate nullspace bases, see [32] for a detailed survey of the available methods.
The remainder of the paper is organized as follows. In Section 2 we present the underlying theoretical foundations of the new schemes. Then, in Section 3 we describe the new schemes. The description of the test problems, the experimental setup and the results are given in Sections 4, 5, and 6, respectively. We conclude and discuss future work in Section 7.
@@ 2 Theoretical Analysis
Consider $A=J+R\in\mathbb{R}^{n\times n}$, where $0\leq R=R^{T}$ and $0\neq J=-J^{T}$. Then there exist a real orthogonal matrix $Q\in\mathbb{R}^{n\times n}$, and integers $n_{1}\geq n_{2}\geq\cdots\geq n_{r-1}>0$ and $n_{r},n_{r+1}\geq 0$, such that
@@
where $0<R_{11}=R_{11}^{T}\in\mathbb{R}^{n_{1}\times n_{1}}$ , $J_{ii}=-J_{ii}^{T}\in\mathbb{R}^{n_{i}\times n_{i}}$ for $i=1,\dots,r$, $J_{rr}$ invertible, and $J_{i,i-1}=-J_{i-1,i}^{T}=[\Sigma_{i,i-1}\;0]\in\mathbb{R}^{n_{i}\times n_{i-1}}$ with $\Sigma_{i,i-1}$ being nonsingular for $i=2,\dots,r-1$.
@@
The proof follows directly from the staircase form of [1] by a further full rank decomposition of the bottom right block.
For a system transformed to the form (6), one can perform a block LDULDU factorization, see [17].
@@
Then it follows that $\hat{C}_{11}^{T}\hat{B}_{11}=K_{1}$ is nonsingular and hence $\hat{C}_{11},\hat{B}_{11}$ are nonsingular, as well as $\hat{B}_{12}=0$, $\hat{C}_{12}=0$. Then the assumption that $\mathop{\rm rank}\nolimits C=\mathop{\rm rank}\nolimits B=k$ implies that $\hat{B}_{22}=0$. Let
@@
Note that the transformation in Theorem 1 is an orthogonal congruence on the $A$ block and hence the property that $A+A^{T}\geq 0$ is preserved under the transformation. Note further that if $B=C$ then $Q_{2}=V_{2}$ and thus $Q=V$ and $B_{21},B_{31},B_{41}=0$.
@@
ii) For the generalized case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B=C$, and $\mathcal{W}$ is invertible, then there exists a real orthogonal matrix $Q$ such that $Q^{T}Z^{T}AZQ=\mbox{\rm diag}(A_{22},A_{33})$, with $A_{33}$ skew-symmetric and invertible. If the matrix $A$ can be transformed to the form (6) with the last two rows and columns void, then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (7), where the last block is void and the symmetric part is positive definite.
iii) For the general case, i.e., if $A\neq A^{T}$, but $A+A^{T}\geq 0$, $B\neq C$, if the matrix $A$ can be transformed to the form (6) with the last two rows and columns void, then the block $A_{33}$ is void and for $A_{22}$ there exists a block $LDU$ factorization as in (7), where the last block is void and the symmetric part is positive definite.
Note that in our algorithmic approaches, the condition in Corollary 1 that $A+A^{T}\geq 0$ can be relaxed that it holds only in the nullspace $Z$.
The construction of the projection matrices $Z,U$ with orthonormal columns will typically lead to full matrices. So for the practical implementation we will use sparse approximate versions of the associated decompositions. In this way we will not obtain exact bases for the nullspaces, but sparse approximations $\tilde{Z},\tilde{U}$. For the symmetric and generalized case, the following result shows that if the error $\|Z-\tilde{Z}\|$ can be bounded by a small quantity, then the approximate nullspace matrix $\tilde{Z}^{T}A\tilde{Z}$ will still have a positive definite symmetric part.
@@
Under the assumptions of Corollary 1, let in the symmetric and generalized case $\tilde{Z}$ be a (sparse) approximation of $Z$ (with columns that are not exactly orthonormal) such that $\|\tilde{Z}-Z\|_{2}\leq\epsilon\|Z\|_{2}=\epsilon$, with a sufficiently small $\epsilon$. Then $\tilde{Z}^{T}A\tilde{Z}$ still has a positive definite symmetric part.
@@ 3 A new multi-layer iterative scheme
where $V\in\mathbb{R}^{n\times n}$ is invertible and $L\in\mathbb{R}^{k\times n}$ is a lower trapezoidal matrix. Then, the rightmost nkn-k columns of $V$ (i.e., $Z=V(:,k+1:n)$ in MATLAB notation) form a basis for the nullspace. Since in our case, we are not interested in $L$, it is not stored. However, since $B$ is sparse, pivoting is usually required to avoid division by zero or small values in modulus. The construction for the nullspace of $C$ is analogous.
A pseudocode for the sparse approximate right oblique conjugation with pivoting to obtain a sparse approximate nullspace basis is presented in Algorithm 2. Hereafter, we refer to this procedure as SAROC. It produces a sparse basis, since we assume that computed coefficients that are smaller than a given threshold and entries of the matrix for which we compute the nullspace that are lower than a given tolerance in absolute value are set to zero, as shown in lines 8 and 9, respectively. As a consequence, we obtain approximate projected nullspace matrices, $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{Z}$ and $\widetilde{\mathcal{N}}=\widetilde{Z}^{T}A\widetilde{U}$, respectively, for the symmetric/generalized and general case.
[…]
4:for $i=1,\ldots,k$ do
5:  $\sigma=b_{i}^{T}[v_{i},v_{i+1},\ldots,v_{n}]$ $\triangleright$ compute coefficient vector $\sigma\in\mathbb{R}^{n-i+1}$
@@
7:  $v_{i}\leftrightarrow v_{i+l^{*}-1}$ $\triangleright$ Swap the corresponding columns of $V$
8:  $\sigma_{1}\leftrightarrow\sigma_{l^{*}}$ $\triangleright$ Swap the corresponding coefficients
@@
Rather than computing the classical Cholesky factorization of the symmetric part $\widetilde{\mathcal{N}}^{s}$ of the approximate projected nullspace matrix, to avoid explicitly forming this matrix, we opt for a factorized approximate inverse based Cholesky factorization. If $\widetilde{\mathcal{N}}^{s}$ is not positive definite, then the Cholesky factorization will detect this. If $\widetilde{\mathcal{N}}^{s}$ is indefinite (with a few negative eigenvalues), it would still be possible to obtain a shifted-skew symmetric system, as shown in [23], but with an additional cost. Alternatively, we could just apply a diagonal shift $\widetilde{\mathcal{N}}^{s}+\alpha I$ but with an additional cost of finding a ’good’ $\alpha$, which can be obtained, for example, by estimating the smallest eigenvalue.
We obtain the FSAI of $\widetilde{\mathcal{N}}^{s}$ using the method in [2]. Our implementation, presented in Algorithm 3, is implicit in the sense that $\widetilde{\mathcal{N}}^{s}$ is not formed explicitly, since it can be relatively dense even if $\widetilde{Z}$, $\widetilde{U}$, and $A$ are sparse. Instead, its columns are extracted, as needed in line 3. This can be easily achieved by implicit matrix-vector multiplications, $n^{s}_{i}=\widetilde{\mathcal{N}}^{s}e_{i}$. Similar to the sparse approximate right oblique conjugation, we replace computed coefficients that are smaller than a given threshold and entries of the matrix that is factorized that are smaller than a given tolerance in absolute value by zero, see lines 5 and 6, respectively. FSAI produces $\widetilde{W}\in\mathbb{R}^{(n-k)\times(n-k)}$ such that $\widetilde{W}^{T}\widetilde{\mathcal{N}}^{s}\widetilde{W}\approx I$. Furthermore, $\widetilde{\mathcal{N}}^{s}$ and also the skew-symmetric part $\widetilde{\mathcal{N}}^{j}$, or any other matrices resulting from further projections on them, do not need to be formed explicitly, since matrix-vector multiplications with them can be computed implicitly.
@@
[…]
4:for $j=1,\ldots,n-k$ do
5:  $\sigma_{j},\sigma_{j+1},\ldots,\sigma_{n-k}]^{T}=[w_{j},w_{j+1},\ldots,w_{n-k}]^{T}n^{s}_{j}$
$\triangleright$ $n^{s}_{j}$ is the $j^{th}$ column of $\widetilde{\mathcal{N}}^{s}$ and $w_{i}$ is the $i^{th}$ column of $\widetilde{W}$
6:  for $i=j+1,\ldots,n-k$ do
7:   if $|\sigma_{i}/\sigma_{j}|>\rho$ then
[…]
@@
[…]